# Domain Generalization in LiDAR Semantic Segmentation Leveraged by Density Discriminative Feature Embedding

### v1

@@ IV Experiments
In this section, we demonstrate the domain generalization performance of our method through extensive experiments. The implementation details and experimental settings of the proposed method are detailed in Sec. IV-A. The configuration of datasets used for evaluation is detailed in Sec. IV-B. A comparative analysis with recent domain generalization and domain adaptation methods is presented in Sec. IV-C. The effectiveness of individual components within the proposed DDFE is analyzed in Sec. IV-D.
@@
| Method | Car | Bicycle | Motor- cycle | Truck | Other vehicle | Pedes- trian | Drivable surface | Side- walk | Walk- able | Vege- tation | mIoU | |
| K | Base | 91.23 | 10.04 | 35.69 | 52.89 | 37.95 | 40.99 | 83.86 | 62.78 | 66.34 | 91.33 | 57.31 |
| DGLSS [10] | 92.76 | 11.99 | 27.09 | 72.50 | 45.95 | 36.39 | 84.76 | 65.64 | 67.98 | 91.28 | 59.62 | |
| Ours | 91.86 | 21.65 | 48.11 | 26.84 | 40.04 | 62.32 | 91.90 | 79.65 | 72.29 | 90.31 | 62.50 | |
| N | Base | 68.91 | 2.51 | 12.18 | 11.30 | 20.35 | 29.47 | 80.17 | 31.91 | 40.19 | 77.18 | 37.42 |
| DGLSS [10] | 76.36 | 1.51 | 35.18 | 26.47 | 25.49 | 37.09 | 82.03 | 38.12 | 44.20 | 81.79 | 44.83 | |
| Ours | 75.29 | 3.84 | 29.28 | 30.11 | 31.37 | 51.69 | 87.85 | 50.02 | 51.10 | 83.77 | 49.43 | |
| W | Base | 72.12 | 2.52 | 4.52 | 7.77 | 13.36 | 40.86 | 64.92 | 30.12 | 34.84 | 81.40 | 35.24 |
| DGLSS [10] | 82.26 | 4.85 | 9.72 | 16.80 | 17.67 | 52.55 | 68.20 | 35.91 | 33.33 | 85.41 | 40.67 | |
| Ours | 83.66 | 4.69 | 13.99 | 25.79 | 18.71 | 61.53 | 63.19 | 38.49 | 28.66 | 88.54 | 42.73 |
IV-A Implementation Details
We employ the point head, inspired by Cylinder3D [1], to produce point-wise outputs as shown in Fig. 3. The process involves duplicating voxel-wise features from the 3D backbone and combining them with the corresponding point-wise features . This amalgamation then undergoes processing through two MLP layers in the point head, culminating in the point-wise segmentation prediction. Simultaneously, the voxel head tackles voxel features with a sparse 3D convolution layer, employing a 333 kernel to create voxel-wise segmentation predictions in line with Cylinder3D’s methodology. For inference, we use only the point head that provides the point-by-point segmentation prediction.
For the loss functions in training, we incorporate the Lovasz-Softmax loss [40] along with the Weighted Cross-Entropy loss . The total loss is calculated by summing the voxel-wise and point-wise losses as follows:
where the weights are uniformly set to 1 for each of the four distinct loss functions. We use the Adam optimizer with an initial learning rate of 1e-3. This rate is decreased by a factor of 0.99 with every epoch. The training is conducted over 30 epochs with a batch size of two on a single RTX 3090 GPU, and the epoch yielding the highest source validation mIoU was chosen. For voxelization, we adopt a cubic size of [20cm, 20cm, 20cm]. In cases where multiple point labels were found within a voxel, the voxel label is determined based on the predominant label, aligning with the approach in Cylinder3D [1].
The total inference time of our method for processing a single frame with the nuScenes dataset on the NVIDIA RTX 3090 is 44ms. It’s important to note that integrating our DDFE module into the MinkNet42 architecture contributes an additional 8ms to the computation time, with the MinkNet42 architecture requiring 36ms. This integration also slightly increases the model complexity, adding about 23.8k parameters (+0.06%). Furthermore, implementing density augmentation during training on the nuScenes dataset requires an additional 60ms per scan.
IV-B Datasets
In our experimental setup, we employ three datasets: Waymo [4], SemanticKITTI [5], and nuScenes [6]. Notably, the Waymo and SemanticKITTI datasets use 64-channel LiDAR, while the nuScenes employs a 32-channel LiDAR. For detailed information, please refer to Fig. 2 and Tab. I. In line with methods from previous studies [12, 10], we split the nuScenes sequences into 700 for training and 150 for validation. The Waymo dataset is partitioned into 798 sequences for training and 202 for validation. Regarding the SemanticKITTI dataset, we use sequences 00 to 10 for training, setting aside sequence 08 exclusively for validation.
To support domain generalization experiments, we standardize the coordinate system across all datasets, aligning it with LiDAR’s coordinate framework. In this unified system, the forward direction of the ego vehicle aligns with the x-axis, the leftward direction with the y-axis, and the upward direction with the z-axis. We also remove the ego vehicle’s point cloud data in the preprocessing phase to enhance computational efficiency. For the LiDAR in the Waymo dataset, it’s noteworthy that the vertical beam inclinations are not uniformly distributed. Detailed information about each vertical beam angle is provided within the dataset itself [4].
| Method | DA | Car | Bicycle | Motor- cycle | Truck | Other vehicle | Pedes- trian | Drivable surface | Side- walk | Terrain | Vege- tation | mIoU | |
| Base | 50.7 | 5.7 | 6.0 | 21.7 | 24.8 | 29.2 | 89.1 | 42.0 | 23.1 | 85.8 | 37.8 | ||
| FADA [41] | ✓ | 69.1 | 1.9 | 21.4 | 42.9 | 20.5 | 12.6 | 84.3 | 43.6 | 38.3 | 80.0 | 41.5 | |
| CLAN [42] | ✓ | 73.1 | 1.7 | 16.3 | 31.4 | 19.1 | 13.5 | 85.7 | 45.7 | 50.1 | 81.4 | 41.8 | |
| DAST [43] | ✓ | 73.2 | 2.1 | 24.4 | 44.6 | 22.4 | 18.2 | 77.6 | 39.5 | 43.2 | 81.2 | 42.6 | |
| DCFNet [8] | ✓ | 74.9 | 1.8 | 28.9 | 33.9 | 21.6 | 19.9 | 84.5 | 44.6 | 46.3 | 83.6 | 44.0 | |
| CutMix [44] | 75.5 | 0.1 | 14.0 | 26.6 | 22.6 | 3.9 | 86.6 | 36.5 | 19.7 | 85.6 | 37.1 | ||
| Copy-Paste [45] | 77.9 | 3.1 | 11.1 | 21.7 | 31.2 | 7.8 | 88.0 | 38.8 | 19.6 | 86.2 | 38.5 | ||
| Mix3D [37] | 72.1 | 0.0 | 34.8 | 11.7 | 26.4 | 28.5 | 83.3 | 41.0 | 46.4 | 86.5 | 43.1 | ||
| PolarMix [46] | 74.1 | 1.7 | 41.9 | 26.9 | 23.8 | 30.5 | 85.1 | 42.7 | 45.3 | 86.2 | 45.8 | ||
| LiDomAug [12] | 79.2 | 5.8 | 28.0 | 49.3 | 32.1 | 13.8 | 88.0 | 42.0 | 35.4 | 85.1 | 45.9 | ||
| Ours | 76.0 | 5.7 | 42.3 | 33.9 | 26.7 | 49.5 | 88.5 | 49.8 | 50.2 | 78.2 | 50.1 | ||
| Base | 78.5 | 0.0 | 8.2 | 3.4 | 11.1 | 34.5 | 66.3 | 35.8 | 39.4 | 84.2 | 36.1 | ||
| FADA [41] | ✓ | 83.3 | 39.0 | 24.5 | 8.7 | 3.6 | 33.6 | 72.7 | 39.4 | 36.0 | 82.1 | 42.3 | |
| CLAN [42] | ✓ | 86.0 | 21.4 | 9.2 | 13.5 | 5.7 | 42.8 | 77.8 | 51.0 | 47.6 | 84.6 | 43.9 | |
| DAST [43] | ✓ | 90.8 | 40.8 | 24.0 | 11.5 | 5.2 | 41.7 | 76.7 | 45.3 | 48.5 | 84.5 | 46.9 | |
| DCFNet [8] | ✓ | 91.7 | 34.3 | 24.5 | 18.5 | 11.6 | 49.1 | 78.6 | 47.5 | 49.8 | 86.4 | 49.2 | |
| CutMix [44] | 81.2 | 0.0 | 5.3 | 9.1 | 17.4 | 11.8 | 73.6 | 45.5 | 46.8 | 85.7 | 37.6 | ||
| Copy-Paste [45] | 85.7 | 0.0 | 8.2 | 12.8 | 6.5 | 28.6 | 80.8 | 47.4 | 53.8 | 87.2 | 41.1 | ||
| Mix3D [37] | 93.1 | 10.4 | 31.3 | 17.0 | 14.1 | 34.2 | 71.8 | 40.7 | 44.6 | 89.5 | 44.7 | ||
| PolarMix [46] | 75.9 | 19.4 | 19.7 | 9.6 | 3.0 | 18.3 | 75.0 | 43.1 | 48.9 | 77.8 | 39.1 | ||
| LiDomAug [12] | 92.6 | 31.6 | 42.5 | 21.6 | 6.4 | 34.4 | 70.0 | 47.1 | 59.4 | 77.5 | 48.3 | ||
| Ours | 89.5 | 4.7 | 29.5 | 22.4 | 6.3 | 33.7 | 81.9 | 50.7 | 56.5 | 88.2 | 46.3 |
| Method | DA | Car | Bicycle | Motor- cycle | Truck | Other vehicle | Pedes- trian | Drivable surface | Side- walk | Terrain | Vege- tation | mIoU | |
| Base | 55.3 | 0.1 | 3.4 | 2.5 | 4.5 | 4.7 | 79.2 | 32.2 | 21.7 | 74.0 | 27.8 | ||
| SWD [47] | ✓ | - | - | - | - | - | - | - | - | - | - | 27.7 | |
| 3DGCA [25] | ✓ | - | - | - | - | - | - | - | - | - | - | 27.4 | |
| C&L [11] | ✓ | - | - | - | - | - | - | - | - | - | - | 31.6 | |
| LiDAR-UDA [30] | ✓ | 73.5 | 0.9 | 15.9 | 0.9 | 25.7 | 40.8 | 87.4 | 42.3 | 47.9 | 83.2 | 41.8 | |
| LiDomAug [12] | - | - | - | - | - | - | - | - | - | - | 39.2 | ||
| Ours | 73.5 | 1.7 | 35.9 | 34.2 | 27.4 | 40.3 | 85.1 | 45.4 | 48.5 | 78.9 | 47.1 | ||
| Base | 58.0 | 0.5 | 4.4 | 3.8 | 5.0 | 12.6 | 33.8 | 2.8 | 30.1 | 80.9 | 23.2 | ||
| SWD [47] | ✓ | - | - | - | - | - | - | - | - | - | - | 24.5 | |
| 3DGCA [25] | ✓ | - | - | - | - | - | - | - | - | - | - | 23.9 | |
| C&L [11] | ✓ | - | - | - | - | - | - | - | - | - | - | 33.7 | |
| LiDAR-UDA [30] | ✓ | 86.2 | 0.0 | 13.9 | 9.3 | 3.2 | 16.5 | 65.7 | 6.1 | 54.1 | 85.7 | 34.0 | |
| LiDomAug [12] | 83.7 | 0.0 | 16.6 | 5.4 | 13.4 | 29.6 | 67.4 | 25.4 | 56.5 | 81.2 | 37.9 | ||
| Ours | 84.6 | 8.7 | 27.6 | 15.8 | 2.6 | 27.6 | 75.6 | 27.6 | 46.8 | 86.1 | 40.3 |
IV-C Comparison to State-of-the-Art DA/DG Methods
The domain of LiDAR-based semantic segmentation generalization is still developing, and standardized experimental protocols are not yet established. Consequently, different methods have employed varied experimental approaches. To ensure equitable comparisons, we adopt the experimental settings used in leading studies such as LiDomAug [12] and DGLSS [10]. In this setup, we compare our method with various Domain Adaptation (DA) techniques. Moreover, we demonstrate the efficacy of our approach by conducting comparisons within the framework of Rangeview-based DA methods, as detailed in [29].
IV-C1 Experiments in the DGLSS Setting
We compare our proposed method with a domain adaptation method COSMIX [7], and domain generalization methods IBN-Net [38], MLDG [39] and DGLSS [10] in the domain generalization experimental setting as introduced in DGLSS [10]. We utilize the Waymo [4], SemanticKITTI [5], and nuScenes [6] datasets. For consistency, we choose MinkowskiNet [24] for our backbone network same as DGLSS.
Results in Tab. II indicate that the proposed method consistently outperforms DGLSS across all datasets, including the Waymo, SemanticKITTI, and nuScenes datasets. Impressively, our method not only demonstrates superior domain generalization performance but also dominates in the source-to-source settings (WW, KK, NN). While DGLSS customizes its augmentation ratios for each dataset, our approach achieves better domain generalization performance using a consistent hyperparameter configuration across all datasets. Our method achieves an average increase of 12.9% over DGLSS for unseen datasets using Waymo as the source data (WK, WN). When SemanticKITTI is the source data, we observe an average improvement of 7.8% for other unseen datasets (KW, KN). Lastly, utilizing nuScenes as the source data results in our approach surpassing DGLSS by an average of 15.8% on other unseen datasets (NW, NK). Tab. III represents the class-specific IoU results using SemanticKITTI as the source dataset. In this context, our method outperforms others in 8 out of 10 classes for the (KN) scenario and in 7 out of 10 classes for the (KW) scenario. These results demonstrate the effectiveness of our method.
| (a) Ground-truth | (b) MinkNet42 | (c) MinkNet42+Ours |
| (a) Ground-truth | (b) MinkNet42 | (c) MinkNet42+Ours |
IV-C2 Experiments in the LiDomAug Setting
We benchmark the proposed method against various augmentation methods, including CutMix [44], Copy-Paste [45], Mix3D [37], PolarMix [46], and domain adaptation methods like SWD [47], 3DGCA [25], Complete & Label [11], LiDAR-UDA [30], CLAN [42], FADA [41], DAST [43] and DCFNet [8]. We employ the experimental setting proposed by LiDomAug [12], using SemanticKITTI [5] and nuScenes [6] datasets. We conduct two distinct experiments: one using the MinkowskiNet [24] as the backbone network and the other using the Complete and Label (C&L) [11].
Tab. IV shows the performance of various data augmentation methods, domain adaptation methods, and our proposed method, all using MinkNet42 as the backbone. The results for CLAN [42], FADA [41], DAST [43], and DCFNet [8] are taken from the DCFNet publication. While our method shows slightly lower performance than LiDomAug in the (NK) scenario, it excels in the (KN) scenario, surpassing all other augmentation methods. Our method demonstrates a significant 9.2% improvement in performance over LiDomAug, which relies on ego-motion and multi-frame information, highlighting its robustness and effectiveness. Moreover, although our approach falls short by 5.9% compared to the latest DA technique, DCFNet [8], in the (NK) scenario, it excels in the (KN) scenario, outperforming DCFNet by 13.9%. The qualitative results, as shown in Fig. 5 and Fig. 6, underscore the significant enhancement in domain generalization performance of the 3D backbone network by our method.
In experiments using C&L [11] as the backbone (see Tab. V), our method notably increases mIoU by 49.1%, 12.7%, and 20.2% in the (KN) scenario compared to C&L [11], LiDAR-UDA [30], and LiDomAug [12], respectively. In the transition from the 32-channel sparse dataset (NK), our approach shows a 19.6% improvement in mIoU over C&L and a 6.3% advantage over LiDomAug. C&L and LiDomAug have an inherent advantage in handling sparse datasets. This advantage is due to their strategy of using multi-frame point clouds combined with known ego-motion to transform sparse data into denser formats. As a result, the difference in performance between these methods and ours is less marked in the (NK) scenario compared to the (KN) context.
| Backbone | Method | Car | Bicycle | Motor- cycle | Other vehicle | Pedes- trian | Truck | Drivable surface | Side- walk | Terrain | Vege- tation | Man- made | mIoU | |
| SalsaNext (Rangeview) | Base | 35.7 | 0.2 | 0.4 | 5.7 | 7.5 | 8.1 | 73.8 | 15.0 | 14.9 | 8.3 | 51.4 | 20.1 | |
| CORAL [48] | 51.0 | 0.9 | 6.0 | 4.0 | 25.9 | 29.9 | 82.6 | 27.1 | 27.0 | 55.3 | 56.7 | 33.3 | ||
| MEnt [49] | 57.4 | 2.2 | 4.6 | 6.4 | 22.6 | 19.3 | 82.3 | 28.8 | 29.9 | 46.8 | 64.2 | 33.1 | ||
| AEnt [49] | 57.4 | 1.1 | 8.6 | 6.7 | 24.0 | 10.1 | 81.0 | 25.4 | 26.6 | 34.2 | 58.9 | 30.4 | ||
| (M+A)Ent [49] | 57.3 | 1.1 | 2.3 | 6.8 | 23.4 | 7.9 | 83.5 | 32.6 | 31.8 | 43.3 | 62.3 | 32.0 | ||
| SWD [47] | 45.3 | 2.1 | 2.2 | 3.4 | 25.9 | 10.6 | 80.7 | 26.5 | 30.1 | 43.9 | 60.2 | 30.1 | ||
| Rochan et al. [29] | 54.4 | 3.0 | 1.9 | 7.6 | 27.7 | 15.8 | 82.2 | 29.6 | 34.0 | 57.9 | 65.7 | 34.5 | ||
| MinkNet42 | Base | 69.7 | 5.3 | 26.2 | 18.3 | 26.5 | 17.5 | 82.2 | 35.6 | 30.6 | 65.2 | 69.7 | 40.6 | |
| (Voxel) | Ours | 75.3 | 1.6 | 44.9 | 15.7 | 28.5 | 19.4 | 87.8 | 48.0 | 44.4 | 67.3 | 72.1 | 45.9 | |
| SalsaNext (Rangeview) | Base | 7.7 | 0.1 | 0.9 | 0.6 | 6.4 | 0.4 | 30.4 | 5.7 | 28.4 | 27.8 | 30.2 | 12.6 | |
| CORAL [48] | 47.3 | 10.4 | 6.9 | 5.1 | 10.8 | 0.7 | 24.8 | 13.8 | 31.7 | 58.8 | 45.5 | 23.2 | ||
| MEnt [49] | 27.1 | 2.0 | 2.3 | 3.4 | 9.5 | 0.4 | 29.3 | 11.3 | 28.0 | 35.8 | 39.0 | 17.1 | ||
| AEnt [49] | 42.4 | 4.5 | 6.9 | 2.8 | 6.7 | 0.7 | 16.1 | 7.0 | 26.1 | 46.1 | 42.0 | 18.3 | ||
| (M+A)Ent [49] | 49.6 | 5.9 | 4.3 | 6.4 | 9.6 | 2.6 | 22.5 | 12.7 | 30.3 | 57.4 | 49.1 | 22.8 | ||
| SWD [47] | 34.2 | 2.7 | 1.5 | 2.0 | 5.3 | 0.9 | 28.8 | 20.5 | 28.3 | 38.2 | 36.7 | 18.1 | ||
| Rochan et al. [29] | 49.6 | 4.6 | 6.3 | 2.0 | 12.5 | 1.8 | 25.2 | 25.2 | 42.3 | 43.4 | 45.3 | 23.5 | ||
| MinkNet42 | Base | 58.7 | 11.3 | 12.6 | 2.1 | 24.2 | 11.6 | 63.3 | 14.7 | 22.1 | 64.5 | 64.9 | 31.8 | |
| (Voxel) | Ours | 87.1 | 15.0 | 20.1 | 4.1 | 20.4 | 16.9 | 83.6 | 45.9 | 54.7 | 81.7 | 76.8 | 46.0 |
IV-C3 Comparison with Rangeview-based DA Methods
To further validate the effectiveness of our method, we perform comparative experiments between our method and various rangeview-based Domain Adaptation (DA) techniques [48, 49, 47], using the 11-class setting from [29]. While these DA methods employ SalsaNext [15], a rangeview-based model, as their baseline, our experiments are based on a 3D voxel-based network, MinkNet42 [24]. The comparative results, except for the MinkNet42 base and our method, are sourced from [29]. Interestingly, even without our proposed method, the MinkNet42 model demonstrates superior generalization performance compared to the rangeview-based backbones with domain adaptation techniques [48, 49, 47, 29]. With the application of our method, MinkNet42’s performance increases by 13.1% in the (KN) scenario and by 44.7% in the (NK) scenario, relative to its baseline.
Rangeview images, trained on uniformly dense 2D projections, often face challenges in domain generalization across different densities. In contrast, 3D voxel-based methods, trained across varying densities in 3D space, are better equipped for adaptation to domain changes. When compared to [29], our method shows significantly higher performance improvements of 33.0% and 95.7% in respective scenarios. The limited performance of the method [29] is likely due to information loss during the rangeview upsampling process, despite their efforts to align projection points between training and target datasets.

### v2

@@ 4 Experiments
In this section, we demonstrate the domain generalization performance of our method through extensive experiments. The implementation details and experimental settings of the proposed method are detailed in Sec. 4.1. The configuration of datasets used for evaluation is detailed in Sec. 4.2. A comparative analysis with recent domain generalization and domain adaptation methods is presented in Sec. 4.3. The effectiveness of individual components within the proposed DDFE and computational cost are analyzed in Section Sec. 4.4. Detailed per-class performance metrics are available in the supplementary material.
4.1 Implementation Details
[…]
We use the Adam optimizer with an initial learning rate of 1e-3. This rate is decreased by a factor of 0.99 with every epoch. The training is conducted over 30 epochs with a batch size of two on a single NVIDIA RTX 3090, and the epoch yielding the highest source validation mIoU is chosen. For voxelization, we adopt a cubic size of [20cm, 20cm, 20cm]. In cases where multiple point labels are found within a voxel, the voxel label is determined based on the predominant label, aligning with the approach in Cylinder3D [48].
4.2 Datasets
In our experimental setup, we employ three datasets: Waymo [35], nuScenes [4], and SemanticKITTI [2]. Notably, the Waymo and SemanticKITTI datasets use 64-channel LiDAR, while the nuScenes employs a 32-channel LiDAR. For detailed information, please refer to Tab. 1. In line with methods from previous studies [29, 16], we split the nuScenes sequences into 700 for training and 150 for validation. The Waymo dataset is partitioned into 798 sequences for training and 202 for validation. Regarding the SemanticKITTI dataset, we use sequences 00 to 10 for training, setting aside sequence 08 exclusively for validation. In integrating class variations across datasets, we follow the class mapping configurations from previous studies for fair comparisons.
4.3 Comparison to State-of-the-Art DA/DG Methods
The field of domain generalization of LiDAR-based semantic segmentation is still in its formative stages, with a lack of universally accepted experimental standards. This situation has led to a diversity of experimental designs across different studies. To facilitate fair comparisons, we align our experimental framework with those utilized in seminal works such as LiDomAug [29] and DGLSS [16]. This approach allows us to benchmark our method against a range of Domain Adaptation (DA) and Domain Generalization (DG) techniques, showcasing the effectiveness of our method in the context of evolving domain generalization challenges.
@@ 4.3 Comparison to State-of-the-Art DA/DG Methods
4.3.1 Experiments in the DGLSS Setting
We compare our method with a domain adaptation method [30] and three domain generalization methods [23, 20, 16], following the experimental framework defined in DGLSS [16] using Waymo (W), SemanticKITTI (K), and nuScenes (N). We adopt MinkowskiNet [7] as our backbone network, aligning with the configuration utilized in DGLSS. The results in Tab. 2 indicate that the proposed method consistently outperforms DGLSS across all datasets. Impressively, our method not only demonstrates superior domain generalization performance but also dominates in the source-to-source settings (WW, KK, NN). Unlike DGLSS, which tailors its augmentation strategies for each specific dataset, our method employs a uniform hyperparameter setting across all datasets, achieving enhanced domain generalization results. Our method achieves an average increase of +12.9% over DGLSS for unseen datasets using Waymo as the source data (WK, WN). With SemanticKITTI as the source, there’s an average enhancement of +7.8% for other unseen domains (KW, KN), and employing nuScenes as the source data yields an average increase of +15.8% on other unseen datasets (NW, NK).
These results demonstrate the robustness of the proposed method in addressing domain discrepancies.
| Backbone | Methods | KN | NK | Backbone | Methods | KN | NK | |
| MinkNet42 [7] | Base | 37.8 | 36.1 | C&L [44] | Base | 27.9 | 23.5 | |
| CutMix [46] | 37.1 | 37.6 | SWD [18] | 27.7 | 24.5 | |||
| Copy-Paste [10] | 38.5 | 41.1 | 3DGCA [39] | 27.4 | 23.9 | |||
| Mix3D [22] | 43.1 | 44.7 | C&L [44] | 31.6 | 33.7 | |||
| PolarMix [40] | 45.8 | 39.1 | LiDomAug [29] | 39.2 | 37.9 | |||
| LiDomAug [29] | 45.9 | 48.3 | LiDAR-UDA [34] | 41.8 | 34.0 | |||
| Ours (v=5cm) | 48.6 | 51.3 | Ours (v=5cm) | 42.5 | 41.0 | |||
| Ours (v=20cm) | 50.1 | 46.3 | Ours (v=20cm) | 47.1 | 40.3 |
4.3.2 Experiments in the LiDomAug Setting
We benchmark the proposed method against various augmentation methods [46, 10, 22, 40], domain adaptation methods [18, 39, 44, 34], and a domain generation method [29]. We adhere to the experimental setup established by LiDomAug [29], employing MinkNet42 [7] and C&L [44] as backbone networks, and utilizing SemanticKITTI and nuScenes datasets for evaluation. Unlike the DGLSS [16] setting which uses a voxel size of 20 cm, the LiDomAug setting uses a voxel size of 5 cm. We benchmark the proposed method against various augmentation methods [46, 10, 22, 40], domain adaptation methods [18, 39, 44, 34], and a domain generation method [29]. We adhere to the experimental setup established by LiDomAug [29], employing MinkNet42 [7] and C&L [44] as backbone networks, and utilizing SemanticKITTI and nuScenes datasets for evaluation. The comparative analysis in Tab. 3 shows that our method surpasses all considered augmentation methods in both (KN) and (NK) configurations. Particularly with MinkNet42 as the backbone, our method achieves a 5.9% and 5.8% increase in performance over LiDomAug in the (KN) and (NK) scenarios, respectively. This enhancement is significant, considering LiDomAug’s reliance on ego-motion and multi-frame data integration for domain generalization. When utilizing C&L [44] as the backbone, the proposed method significantly improves mIoU by +34.5% and +8.4% for the scenario of training on SemanticKITTI and evaluating on nuScenes (KN), compared to C&L and LiDomAug, respectively. Moreover, in the transition from the sparse 32-channel dataset (nuScenes) to the denser 64-channel dataset (SemanticKITTI) (NK), our method shows a +21.7% increase in mIoU over C&L and +8.2% improvement over LiDomAug.
NK
KN
We further conduct an analysis with a voxel size of 20 cm, unlike the comparative methods that all use a voxel size of 5 cm. Increasing the voxel size from 5 cm to 20 cm results in reductions of 30.3% in training time and 62.5% in inference time, while maintaining comparable performance—with a slight decrease of 3.5% on the MinkNet42 backbone and an increase of 4.7% on the C&L backbone. Based on these results, we recommend a voxel size of 20 cm (v=20cm) as the default setting for our proposed method. The effectiveness is further illustrated in the qualitative results depicted in Fig. 3, highlighting the substantial improvements our method brings to the domain generalization performance of the 3D backbone network. Please refer to the supplementary material for detailed results, including per-class IoU.